Consider an incumbent that is a monopoly currently earning $1 million annually. Given the declining costs of raw materials, the incumbent believes a new firm may enter the market. If successful, a new entrant would reduce the incumbent's profits to $750,000 annually. To keep potential entrants out of the market, the incumbent lowers its price to the point where it is earning $850,000 annually for the indefinite future. If the interest rate is 5 percent, does it make sense for the incumbent to limit price to prevent entry?

Yes, since $250,000 > $5 million.

Yes, since $2 million > $150,000.

No, since $5 million > $100,000.

Yes, since $2 million > $250,000.
Answer #1

Profit from setting the limit price = 850000-750000= 100000

At an interest rate of 5%,the profit for an indefinite period = 100000/0.05 = 2000000

Loss from not setting the limit price = 100000-750000 = 250000

Since the loss of 250000 < profit of 2000000 so the firm should set the limit price

option(D)